OpenStudy (anonymous):

determine the slope with the numbers (0,-5) and (-4,0)

OpenStudy (anonymous):

-5/4

OpenStudy (anonymous):

(y2-y1)/(x2-x1) = [-5-0]/[0-(-4)] = -5/4